# RNAGenScape: Property-guided Optimization and Interpolation of mRNA Sequences with Manifold Langevin Dynamics

### v1

@@ 7 Related Works
De novo generative models excel at creating novel sequences, but fundamentally operate by generating from scratch rather than refining existing functional sequences (Prykhodko et al., 2019; Méndez-Lucio et al., 2020; Dauparas et al., 2022; Wu et al., 2021; Madani et al., 2023; Watson et al., 2023). Classic optimization strategies (Williams, 1992; Zinkevich, 2003; Brooks, 1998; Andrieu et al., 2003; Selman & Gomes, 2006) are capable of improving known sequences, but typically lack mechanisms to ensure that intermediate variants remain consistent with the underlying biological distribution. More recent deep learning methods for sequence generation and optimization (Luo et al., 2022; Shuai et al., 2023; Gruver et al., 2023) aim to combine generative modeling with property-driven objectives, but their optimization trajectories remain opaque, offering limited interpretability of which sequence changes drive functional improvements. Furthermore, existing methods cannot interpolate between sequences. Consequently, a framework that enables biologically-grounded sequence engineering remains needed (Wu et al., 2019).
An extended related works section can be found in Appendix E.
@@ Appendix A Hyperparameters and Architecture
Zebrafish 5.0 / 1.0 {1.0, 0.8, 0.5} $1\times 10^{-2},\,8\times 10^{-3}$
@@
Ribosome-loading 1.0 / 1.0 {0.3} $5\times 10^{-3},\,1\times 10^{-2}$
@@ Appendix B Evaluation Metrics
To quantify the effectiveness of property optimization of different models, we measure both the median improvement and the fraction of sequences that are successfully optimized.
@@
Here, $\Delta_{\text{median}}$ measures the improvement in the target property across the test set, while $\%_{\text{success}}$ reports the percentage of sequences that improve after optimization.
@@
Manifold Fidelity
Given a property prediction model $\mathcal{P}_{\text{oracle}}$ with an encoder $\mathcal{E}_{\text{oracle}}$, a test set of mRNA sequences $X_{\text{test}}$, and generated or optimized sequences $\tilde{X}_{\text{test}}$, we quantify manifold fidelity as the average minimum $\ell_{2}$ distance in the latent space of $\mathcal{E}_{\text{oracle}}$ between each new sample and the test data:
$\mathcal{M}_{\text{fidelity}}=\mathbb{E}_{\tilde{x}\sim\tilde{X}_{\text{test}}}\left[\min_{x\in X_{\text{test}}}\big\|\mathcal{E}_{\text{oracle}}(\tilde{x})-\mathcal{E}_{\text{oracle}}(x)\big\|_{2}\right].$ (7)
This metric captures how closely the generated or optimized sequences remain to the empirical data manifold defined by $X_{\text{test}}$.
Appendix C Ablation on SUGAR
We performed ablation on the upsampling ration in SUGAR and the results are summarized in Table S2. In theory, SUGAR is most helpful in datasets that show greater sparsity and more undersampled regions on the data manifold.
In our ablation studies, we observe that different datasets require different ideal SUGAR ratios. In general, smaller datasets need higher upsampling to achieve the optimal performance. While dataset size not necessarily reflect the density or sparsity of the data manifold, in general they seem to be positively correlated. In future practice, we suggest using higher SUGAR upsampling ratio when working with smaller datasets, which is very common in the world of mRNA design.
Table S2: Ablation on SUGAR.
@@ Appendix C Ablation on SUGAR
++property -property ++property -property ++property -property
SUGAR ratio $\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$ $\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$ $\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$
@@
0.5 -0.51 70.3 -0.89 75.1 -0.13 58.1 -0.38 72.5 0.21 58.5 -1.33 81.0
[…]
Appendix D Additional optimization trajectories
Figure S1: More examples of latent space trajectories. (a) Trajectories in the PHATE space. (b) 2D structures. (c) 3D structures.
Figure S2: More examples of latent space trajectories. (a) Trajectories in the PHATE space. (b) 2D structures. (c) 3D structures.
Figure S3: More examples of latent space trajectories. (a) Trajectories in the PHATE space. (b) 2D structures. (c) 3D structures.
Appendix E Extended Related Works
@@
Generative models enable sampling of novel sequences enriched for desired traits. Variational autoencoders (VAEs) (Kingma et al., 2013) have been applied to proteins to learn smooth latent spaces that are amenable to gradient-based optimization (Sinai et al., 2017; Castillo-Hair et al., 2024). ProteinMPNN (Dauparas et al., 2022), although described as a message-passing neural network by the authors, shares core design principles with autoencoders. Generative adversarial networks (Goodfellow et al., 2020) such as Méndez-Lucio et al. (Méndez-Lucio et al., 2020) or ProteinGAN (Wu et al., 2021) and autoregressive language models such as ProGen (Madani et al., 2023) have also been used to generate diverse protein sequences. More recently, diffusion models (Ho et al., 2020) have shown promise in discrete domains. For example, RFdiffusion (Watson et al., 2023) generates proteins unconditionally or conditioned on structural constraints. These methods can be readily adapted to mRNA design.
@@
Sequence optimization can be framed as a black-box search or a differentiable surrogate-guided process. Several approaches relax discrete inputs for gradient-based updates, such as using straight-through estimators (Linder et al., 2019). ReLSO learns a continuous latent space and performs gradient ascent (Castro et al., 2022). Others apply reinforcement learning (Eastman et al., 2018) or Monte Carlo algorithm (Wirecki et al., 2023) for sequence optimization. Methods such as Fast SeqProp (Linder & Seelig, 2021) and LaMBO (Stanton et al., 2022) have demonstrated success in optimizing sequences under multi-objective constraints.

### v2

@@ 5 Related Works
De novo generative models excel at creating novel sequences, but fundamentally operate by generating from scratch rather than refining existing functional sequences (Prykhodko et al., 2019; Méndez-Lucio et al., 2020; Dauparas et al., 2022; Wu et al., 2021; Madani et al., 2023; Watson et al., 2023). Classic optimization strategies (Williams, 1992; Zinkevich, 2003; Brooks, 1998; Andrieu et al., 2003; Selman and Gomes, 2006) are capable of improving known sequences, but typically lack mechanisms to ensure that intermediate variants remain consistent with the underlying biological distribution. More recent deep learning methods for sequence generation and optimization (Luo et al., 2022; Shuai et al., 2023; Gruver et al., 2023) aim to combine generative modeling with property-driven objectives, but their optimization trajectories remain opaque, offering limited interpretability of which sequence changes drive functional improvements. Furthermore, existing methods cannot interpolate between sequences. Consequently, a framework that enables biologically-grounded sequence engineering remains needed (Wu et al., 2019).
An extended related works section can be found in Supplementary Section F.
6 Conclusion
We introduced RNAGenScape, a manifold Langevin dynamics framework for property-optimized mRNA sequence generation. Rather than optimizing in the ambient sequence space, RNAGenScape operates directly on a learned manifold of viable mRNA sequences. In this setting, property gradients are more reliable, biological viability is better preserved, and optimization can be performed with lower computational overhead. Comprehensive experiments on three real-world mRNA datasets show that our method leads to greater property gains and higher success rates under realistic data regimes. With this work, we aim to shift the paradigm of biological sequence design from unconstrained generation toward guided optimization, and to highlight mRNA sequence design as a critical yet understudied frontier in computational biology.
7 Limitations and Future Work
One limitation of RNAGenScape is that it exclusively models and optimizes mRNA sequences only, without explicitly incorporating mRNA structure. While sequence-based representations capture many aspects of mRNA function, structures can provide auxiliary information in regulating translation, stability, and interactions.
In future work, we will study the possibility of performing sequence-structure joint modeling and optimization. Beyond mRNA, we plan to extend RNAGenScape to other modalities such as protein sequences and regulatory elements, and integrate active learning frameworks that guide wet lab experimentation. By grounding sequence optimization in the manifold of real data, we aim to provide a versatile platform for interpretable high-throughput design in synthetic biology.
Impact Statement
This work proposes a machine learning framework for optimizing mRNA sequences while preserving biological viability by constraining optimization to a learned data manifold. Potential positive impacts include supporting computational design workflows in mRNA therapeutics and synthetic biology such as vaccine development by enabling more controlled and interpretable sequence refinement. As a general machine learning contribution, the framework may also inform broader research on manifold-constrained optimization for biological sequences.
The approach is intended as a computational tool to complement experimental and domain expertise rather than a substitute for empirical validation. As with other data-driven sequence models, limitations and biases in the available training data may influence optimization outcomes, and inappropriate use without biological oversight could lead to nonfunctional or misleading designs.
We do not anticipate direct negative societal consequences arising from this work beyond those already associated with machine learning-based biological modeling. We encourage future research to further study robustness, data bias, and safe integration of such methods into experimental pipelines.
@@ Appendix A Hyperparameters and Architecture
Zebrafish 5.0 / 1.0 {1.0, 0.8, 0.5} $1\times 10^{-4},\,8\times 2^{-4}$
@@
Ribosome 1.0 / 1.0 {0.3} $5\times 10^{-3},\,1\times 4^{-3}$
@@ Appendix B Evaluation Metrics
To quantify the effectiveness of property optimization of different models, we measure both the median property gain and the fraction of sequences that are successfully optimized.
@@
Here, $\Delta_{\text{median}}$ measures the gain in the target property across the test set, while $\%_{\text{success}}$ reports the percentage of sequences that improve after optimization.
@@
The external oracle
In our evaluation for property optimization, we use an external oracle $\mathcal{P}_{\text{oracle}}(x)$ to predict the property of sequences generated by RNAGenScape as well as competing methods. Specifically, we adopt UTR-LM (Chu et al., 2024), a widely used mRNA language model, as the property prediction oracle. It is first pre-trained on more than 480,000 5’ UTR mRNA sequences from diverse sources, and then fine-tuned for property prediction on each of our datasets. After fine-tuning, it is frozen and never accessed during optimization, ensuring unbiased evaluation. Performance of the fine-tuned UTR-LM is shown in Table S2.
Table S2: The UTR-LM oracle for property evaluation.
Property Prediction
Pearson Corr. $\uparrow$ Spearman Corr. $\uparrow$
@@
Appendix C Property optimization performance against all baselines
Figure S1: Quantitative comparison on property optimization. Median change in property (left) and success rate (right) are color-coded such that darker results are better. Properties of generated mRNA sequences are evaluated with the UTR-LM oracle pre-trained on more than 480,000 5’ UTR mRNA sequences and fine-tuned on each of the three datasets. OpenVaccine ($n\approx 2k$, property: stability), Zebrafish ($n\approx 55k$, property: translation efficiency), and Ribosome ($n\approx 260k$, property: mean ribosome load).
Appendix D Interpolation between sequences
In addition to optimizing mRNA sequences, we can interpolate between two existing sequences by guiding the latent embedding of one sequence toward that of another. Specifically, given a source sequence $x_{\text{source}}$ and a target sequence $x_{\text{target}}$, we first obtain their latent embeddings via the encoder: $z_{\text{source}}=\mathcal{E}(x_{\text{source}})$ and $z_{\text{target}}=\mathcal{E}(x_{\text{target}})$.
We then perform property-guided Manifold Langevin dynamics starting from $z=z_{\text{source}}$, with the drift term being the force term that drives the embedding toward $z_{\text{target}}$:
$f_{\text{interp}}(z,z_{\text{target}})=-\frac{z-z_{\text{target}}}{\|z-z_{\text{target}}\|_{2}}$ (6)
In this case, we slightly modify the update rule of the Langevin dynamics (equation 3) as follows:
$\text{d}z_{t}=\eta\left(\nabla_{z}f(z_{t})+\lambda f_{\text{interp}}\right)+\sqrt{2\eta}\cdot\epsilon_{t}.$ (7)
By setting the interpolation weight $\lambda>0$ in equation 7, we add a directional bias that drives the latent trajectory toward the target point, while following the property gradients.
Results on interpolation
Guided by a directional force toward a specified target, RNAGenScape generates smooth and coherent trajectories on the learned manifold while preserving biological plausibility and continuity (Figure S2). These trajectories connect arbitrary input-target sequence pairs in a structured manner, reflecting semantically meaningful transitions. The distances from each intermediate point to the source and target quantitatively demonstrate the monotonicity and smoothness of the interpolation (Figure S3).
Figure S2: Latent space interpolation trajectories from 5 sources to 4 targets. Each trajectory is shown as a line fading from bright to dark in a consistent color. RNAGenScape produces smooth and coherent paths on the manifold between arbitrary input-target mRNA pairs.
Figure S3: Latent space $\ell_{2}$ distances during interpolation show smooth and monotonic transition from the source to the target. Results are averaged over all data samples.
Appendix E Additional optimization trajectories
Figure S4: More examples of latent manifold trajectories. Trajectories in PHATE space and corresponding 3D structures are shown.
Figure S5: More examples of latent manifold trajectories. Trajectories in PHATE space and corresponding 3D structures are shown.
Figure S6: More examples of latent manifold trajectories. Trajectories in PHATE space and corresponding 3D structures are shown.
Appendix F Extended Related Works
@@
Generative models enable sampling of novel sequences enriched for desired traits. Variational autoencoders (VAEs) (Kingma et al., 2013) have been applied to proteins to learn smooth latent manifolds that are amenable to gradient-based optimization (Sinai et al., 2017; Castillo-Hair et al., 2024). ProteinMPNN (Dauparas et al., 2022), although described as a message-passing neural network by the authors, shares core design principles with autoencoders. Generative adversarial networks (Goodfellow et al., 2020) such as Méndez-Lucio et al. (Méndez-Lucio et al., 2020) or ProteinGAN (Wu et al., 2021) and autoregressive language models such as ProGen (Madani et al., 2023) have also been used to generate diverse protein sequences. More recently, diffusion models (Ho et al., 2020) have shown promise in discrete domains. For example, RFdiffusion (Watson et al., 2023) generates proteins unconditionally or conditioned on structural constraints. These methods can be readily adapted to mRNA design.
@@
Sequence optimization can be framed as a black-box search or a differentiable surrogate-guided process. Several approaches relax discrete inputs for gradient-based updates, such as using straight-through estimators (Linder et al., 2019). ReLSO learns a continuous latent manifold and performs gradient ascent (Castro et al., 2022). Others apply reinforcement learning (Eastman et al., 2018) or Monte Carlo algorithm (Wirecki et al., 2023) for sequence optimization. Methods such as Fast SeqProp (Linder and Seelig, 2021) and LaMBO (Stanton et al., 2022) have demonstrated success in optimizing sequences under multi-objective constraints.